OpenStudy (anonymous):

How can one sixthx − 5 = one fifthx + 2 be set up as a system of equations? 6y − x = −30 5y − x = 10 5y + x = −30 5y + x = 10 6y − 6x = −30 5y − 5x = 10 5y + 5x = −30 5y + 5x = 10

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@mathmate @mathmale @Hotchellerae21 @dan815

OpenStudy (mathmate):

You would divide each equation by the coefficient of y. Then isolate y to the form y=a1x+b1, and y=a2x+b2 Equate the two right-hand sides to give one single equation in x, from which x can be found.

OpenStudy (anonymous):

ok so it would be C

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

it would be C

OpenStudy (mathmate):

sorry, C is not the correct answer.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so what would it bee?

OpenStudy (mathmate):

If you take the first equation in A, 6y − x = −30 and divide by 6, you'd get y-x/6=-5, or y=x/6-5 which is the first part of the given equality. Check the other equation to see if it works out to be the other part of the given equation.

OpenStudy (anonymous):

ok so it would be A

OpenStudy (mathmate):

I hope you worked out the second equation, because I didn't.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

@mathmate i did

OpenStudy (mathmale):

Andy, please explain your reasoning before you type in "so it would be A." It's how you obtain your answer, or at least how you try to obtain it, that is important for long term understanding and recollection. You are given "How can one sixthx − 5 = one fifthx + 2 be set up as a system of equations?" and you are asked to take this ONE equation involving the variable x only and to re-write it as two separate equations in y and x both; you are not asked to actually solve this system of equations. Take the given equation: "one sixthx − 5 = one fifthx + 2" and separate it at the = sign into two halves: the left half and the right half; label each part with " y= ". Try that now, please.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

If you take the first equation in A, 6y − x = −30 and divide by 6, you'd get y-x/6=-5, or y=x/6-5 which is the first part of the given equality. @mathmale

OpenStudy (anonymous):

Its going to be A i can feel it.

OpenStudy (anonymous):

@geekfromthefutur will it be A because it think its going to be A because the first part of the equation works and thats the only answer that uses the first equation so i think its A. @mathmale

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

can you help me @campbell_st

OpenStudy (anonymous):

ill give you a medal

OpenStudy (campbell_st):

well it looks like people have given you some help

OpenStudy (anonymous):